OpenStudy (anonymous):

Evaluate x over y, for x = 2/3 and y = 5/6

OpenStudy (anonymous):

a. 5/9 b. 4/5 c. 5/4 d. 9/5

OpenStudy (anonymous):

\[\frac{2}{3}\div\frac{5}{6}=\frac{2}{3}\times \frac{6}{5}\] is a start

OpenStudy (anonymous):

b?

OpenStudy (anonymous):

correct